OpenStudy (anonymous):

Thirteen chickens and pigs are at the backyard feeder. You count 36 legs. How many chickens and how many pigs are there?

OpenStudy (anonymous):

13 * 2 = 26. 36 - 26 = 10 10 / 4 = 2.5 ....hmmm, did I mess up somewhere? Oh well there are 2 pigs. and a messed up chicken

OpenStudy (anonymous):

5 pigs and 8 chickens 5(4)+8(2)=36

OpenStudy (jamesj):

The way to solve it is something like this. Let C = # of chickens P = # of pigs Then C + P = 13 Now as 2C = # of chicken legs and 4P = # pig legs, we have 2C + 4P = 36. Hence we have two equations in C and P: C + P = 13 2C + 4P = 36 Now solve that however you normally solve such equations.